# Scaling Intelligence: Designing Data Centers for Next-Gen Language Models

### v1

@@ Abstract.
The explosive growth of Large Language Models (LLMs)—such as GPT-4 with 1.8 trillion parameters—demands a radical rethinking of data center architecture to ensure scalability, efficiency, and cost-effectiveness. Our work provides a comprehensive co-design framework that jointly explores FLOPS, HBM bandwidth and capacity, multiple network topologies (two-tier vs. FullFlat optical), the size of scale-out domain, and popular parallelism/optimization strategies used in LLMs. We introduce and evaluate Fullflat network architectures, which provide uniform high-bandwidth, low-latency connectivity between all nodes, and demonstrate their transformative impact on performance and scalability. Through detailed sensitivity analyses, we quantify the benefits of overlapping compute and communication, leveraging hardware-accelerated collectives, widen scale-out domain and larger memory capacity. Our study spans both sparse (mixture of experts) and dense transformer-based LLMs, revealing how system design choices affect Model FLOPS Utilization (MFU = Model flops per token $\times$ Observed tokens per sec/Peak flops of the hardware) and overall throughput. For co-design study, we extended and validated a performance modeling tool capable of predicting LLM runtime within 10% of real-world measurements. Our findings offer actionable insights and a practical roadmap for designing AI data centers that can efficiently support trillion-parameter models, reduce optimization complexity, and sustain the rapid evolution of AI capabilities.
@@ 2.2.1. Model Components:
[…]
@@ 2.3. Systems and Their Parameters
Prior research (Yin et al., 2013) and recent advancements (Patel et al., 2025; Lightmatter, 2025; Systems, 2025) suggest the feasibility of a fully optical networks with co-packaged optics (CPO) for both HBD/LBD domains, offering high bandwidth, reduced power consumption, and cost-effectiveness. These networks can adopt 2D or 3D configurations, such as HyperX or PolarFly, allowing access to any GPU in two or three hops. A Full flat network use advanced CPO to equalize bandwidth between HBD and LBD, though communication outside HBD incurs higher latency due to multiple hops. Figure 4 shows a few potential Full flat topologies, with GPUs linked via single-hop connections within HBD, while those requiring multiple switches are part of LBD. FullFlat employs 2D/3D HyperX, Polar* or similar meshes, ensuring each node is $\leq$3 hops from any other at equal HBW per hop. Adaptive routing mitigates “traffic storms” and reduces tail latency. FullFlat equalizes HBD and LBD (high and low bandwidth domains), allowing all traffic types on a single network without reconfiguration. This removes the scale-up/scale-out partitioning required in two-tier fabrics, simplifying scheduling and model mapping. Full flat refers to full bisection bandwidth over a folded high-radix, multi-stage optical mesh, not $N^{2}$ physical links. FullFlat offers adaptive routing, flexible workload placement, and unique parallelization opportunities. It is less sensitive to missing optimizations (compute-comm overlap, hardware collectives), with only a 5% performance gap across top 5000 configs (vs. 80% for two-tier), which should translate to less performance engineering effort (1). Additionally, it should also improve resiliency, robustness, and require less redundancy tax since any GPU can replace any faulty one anywhere in the network. While technical details of the Full flat network will likely be disclosed in a future paper, this study anticipates the development of such networks by 2026/2027.
@@
Table 3 outlines the system specifications used in our co-design analysis, informed by projections of future technological landscapes. The configuration FullFlat refers to the full flat system with equal scale-up and scale-out bandwidths. TwoTier-HBD64—128 represents upcoming two-tiered network with around 8×8$\times$8 × difference between scale-up and scale-out bandwidths, experimented with HBD sizes of 64 and 128. TwoTier-HBD8 reflects current systems with an HBD size of 8. The GPUs used for FullFlat and TwoTier-HBD64—128 are of same computational power and storage, whereas GPUs used in TwoTier-HBD8 are weaker. Our co-design study examines LLM performance sensitivity to various system parameters in Table 3, providing insights into the system requirements for such models. We considered two-tiered networks with scale-up bandwidths of 450-3600 GB/s and scale-out bandwidths of 50-1600 GB/s, covering existing and potential future systems. The FLOPS FP8 and FP16 rows indicate floating-point operations per second for FP8 and FP16 precision, measured in petaflops. Tier 1 and Tier 2 Memory Bandwidth represent data transfer rates within the memory hierarchy, with Tier 1 denoting faster memory like HBM and Tier 2 being standard DDR/LPDDR memory. Memory Capacity reflects storage capacity in each tier. Scale-up and scale-out Bandwidth values describe intra-node/HBD and inter-node/LBD bandwidth capabilities. Latency indicates the minimum delay in data transfer for each network tier. Cluster Size denotes the number of nodes within each system. We anticipate systems of this scale will be available by 2026, and we analyze their performance on large-scale LLMs.
@@ 3. Co-design Study
[…]
@@ 3.1. Strong Scaling with Cluster Size
By 2026, systems are anticipated to run the GPT-1.8T model at speeds $50$-70×70$\times$70 × faster (at 4K gpus, TwoTier-HBD8 vs others), assuming advanced networking infrastructures are available. However, increasing the number of GPUs may lead to performance degradation beyond a certain point due to overheads compared to the pure compute being done. This underscores the importance of strategically investing in advanced network architectures for large-scale deployments of LLMs.
@@ 3.2. Compute/Communication Overlap
High-performance interconnects like NVIDIA Mellanox InfiniBand and Intel Tofino switches (Corporation, 2020) support hardware-accelerated in-network collectives for scale-out domains, while NVSwitches offer accelerated collectives for scale-up domains. Nvidia’s SHARP FP8 technology allows CPUs and GPUs to offload collective tasks to the network chip (Moore, 2025; Graham et al., 2020) improving performance. The analytical model considers traffic volume, latency reduction, and GPU cycle savings (about 13%) due to the presence of a hardware-supported collective engine. Figure 5 (c) illustrates the performance slowdown from using software-based collectives, with 2×2$\times$2 × traffic for all-reduce and 1.5×1.5$\times$1.5 × for reduce-scatter and all-gather. The y-axis shows the percentage slowdown, with the number of GPUs varying from 64 (512 for GPT-29T) to 65,536 shown in the x-axis.
@@ 3.5. Sensitivity to Scale up (SU) Bandwidth
The most significant performance boost occurs when the SU bandwidth doubles from 450 GB/s to 900 GB/s. For GPT-1.8T with an HBD of 128, a 2×2$\times$2 × increase in SU bandwidth results in a 1.6×1.6$\times$1.6 × performance improvement. In contrast, with an HBD of 64, the improvement is 1.4×1.4$\times$1.4 × because it relies on the slower SO network for some expert communications when using 8K GPUs. When the HBD is set to 128, experts can fully utilize the high-speed HBD/SU domain, leading to over a 2.62×2.62$\times$2.62 × improvement from an 8×8$\times$8 × increase in SU bandwidth. For GPT-29T, the performance is similar for both HBD64 and HBD128, as expert communication primarily depends on the SO domain in both scenarios. In this case, the 8×8$\times$8 × increase in SU bandwidth provides a 1.9×1.9$\times$1.9 × improvement.
@@ 3.6. Sensitivity to Scale-Out (SO) Bandwidth
Figure 5 (f) shows the impact of doubling scale-out (SO) bandwidth on GPT4 MOE models’ performance, with scale-up bandwidth fixed at 1,600 GB/s at 80% efficiency. We tested sensitivity for HBD sizes 64 and 128. For GPT-1.8T, in the HBD64 scenario, expert parallelization and communication occur through the scale-out domain, leading to consistent performance improvements with increased SO bandwidth (1.36×1.36$\times$1.36 × at 3.6TB/s). In the HBD128 case, since experts fit within HBD, improvements are limited to data and pipeline parallel communications with minimal impact (1%) on overall performance. For GPT-29T, in both HBD64 and HBD128 scenarios, most expert communications occur through the scale-out domain, resulting in steady performance gains with increased SO bandwidth (1.35×1.35$\times$1.35 × at 3.6TB/s). HBD128 shows slightly more improvements than HBD64 due to faster communication (i.e., all-to-all) within the HBD domain that the experts can leverage.
@@ 3.7. Sensitivity to Compute FLOPS
Figure 5 (g) illustrates the sensitivity of GPT4 MOE’s performance to varying FLOPS per GPU. The SU bandwidth is fixed at 1.6 TB/s, and the SO bandwidth is set at 200 GB/s for the TwoTier-HBD* configuration. The relative speedup compared to the baseline: TwoTier-HBD64 at 2.3 FP16 TFLOP/s per GPU. As FLOPS increase from 2.3 PF/s to 18.4 PF/s, throughput improves across all configurations: TwoTier-HBD64, TwoTier-HBD128, and FullFlat. For the GPT-1.8T model, TwoTier-HBD64 sees a 1.4×1.4$\times$1.4 × boost, while TwoTier-HBD128 and FullFlat improve by 1.59×1.59$\times$1.59 × and 1.66×1.$6$, respectively. The performance gap between TwoTier-HBD64 and TwoTier-HBD128 ranges from 25% to 39%. For the GPT-29T model, TwoTier-HBD64 and TwoTier-HBD128 exhibit performance gains around 1.62×1.62$\times$1.62 ×, whereas FullFlat achieves a 1.9×1.9$\times$1.9 × gain with increased FLOPS.
@@ 3.8. Sensitivity to Memory Bandwidth
Figure 5 (h) shows the impact of varying HBM bandwidth from 3 TB/s to 48 TB/s across three configurations: TwoTier-HBD64, TwoTier-HBD128, and FullFlat, evaluated on 8,192 GPUs. For GPT-1.8T, increasing HBM bandwidth results in a 4.5×4.5$\times$4.5 × performance improvement with a 16×16$\times$16 × bandwidth increase. FullFlat consistently delivers the highest throughput, followed closely by TwoTier-HBD128, due to its capacity to accommodate most communications within the scale-up (HBD=128) domain for GPT-1.8T, resulting in less than a 1% performance gap between TwoTier-HBD128 and FullFlat.
For GPT-29T, increasing HBM bandwidth enhances performance across all configurations, achieving up to 3.2×3.2$\times$3.2 × improvement. FullFlat consistently gives the best performance, while TwoTier-HBD64 and TwoTier-HBD128 exhibit similar performance because both HBD domains cannot accommodate all TP/ES/EP communications in the SU domain, necessitating substantial use of the slower scale-out domain.
@@ 3.9.1. Impact of Tier-1 Memory:
The benefit of co-design study is that it tells us exactly when to stop adding more resources in the system as it won’t bring provide any more ROI. This analysis indicates that having a 1.3TB capacity is a decent target for data centers for the next couple of years. Increasing HBM capacity boosts throughput significantly (4.9×4.9$\times$4.9 × for GPT-1.8T). Thus, optimizing HBM capacity alongside other system parameters is crucial for achieving optimal performance for LLMs.
@@ 3.10. Non MOE Dense LLM: GPT3-175B
Figure 7 shows the GPT-3 175B model’s performance sensitivity across different system setups. Throughput can be 10×10$\times$10 × higher, with optimal scaling stabilizing at 32K GPUs. Performance is compute-bound until communication and other overheads become significant. Overlapping compute and communication is vital for smaller models, as neglecting this can cut performance by 43% at 16,384 nodes. Lack of hardware collectives may lead to a 29% drop. While HBD = 64 suffices for GPT-3, HBD = 128 is optimal for all models analyzed, given a scale-out bandwidth of 200 GB/s or more. This underscores the necessity for fast, efficient networks, hardware-accelerated collectives, and topology-aware communication libraries to effectively support LLM models in future data centers.
@@
Average: TwoTier-HBD8 6%, TwoTier-HBD64 6%, FullFlat 4%
@@
Table 6 presents the percentage of exposed communication and overhead times (recomputation, pipeline bubble, exposed offload times) for the three models discussed. MoE models exhibit significantly higher communication and overhead times compared to smaller dense models. The dense GPT3 model has higher arithmetic intensity and lower memory and network intensities compared to the MOE models. As a result, compared to the sparse MOE models, the dense model demonstrates a lower sensitivity to changes in scale-up and scale-out bandwidth. Sensitivity to HBM bandwidth increase is also lower (less than 3×3$\times$3 × vs 4.5×4.5$\times$4.5 × for the MOE model). However, flop sensitivity is higher than MOE counterparts (3×3$\times$3 × vs 2×2$\times$2 ×).
@@
#GPUs (128$\times$) 89$\times$ perf gain 11$\times$ perf gain from 16$\times$ more GPUs 100$\times$ perf gain
@@
FLOPS (8$\times$) 1.66$\times$ perf gain 1.93$\times$ perf gain 2.73$\times$ perf
HBM BW (16$\times$) 4.2$\times$ perf gain 3.24$\times$ perf gain 2.63$\times$ perf gain
HBM Cap (2TB/GPU) 3.75$\times$ perf gain 1.2$\times$ perf gain N/A
@@
Scale UP BW (8$\times$) 2.62$\times$ perf gain 1.9$\times$ perf gain 1.59$\times$ perf gain
Scale Out BW (8$\times$) 1.36$\times$ perf gain 1.37$\times$ perf gain 1.13$\times$ perf gain

### v2

@@ Abstract.
The explosive growth of Large Language Models (LLMs)—such as GPT-4 with 1.8 trillion parameters—demands a fundamental rethinking of data center architecture to ensure scalability, efficiency, and cost-effectiveness. Our work provides a comprehensive co-design framework that jointly explores FLOPS, HBM bandwidth and capacity, multiple network topologies (two-tier vs. FullFlat optical), the size of scale-out domain, and popular parallelism/optimization strategies used in LLMs. We introduce and evaluate Fullflat network architectures, which provide uniform high-bandwidth, low-latency connectivity between all nodes, and demonstrate their transformative impact on performance and scalability. Through detailed sensitivity analyses, we quantify the benefits of overlapping compute and communication, leveraging hardware-accelerated collectives, widening the scale-out domain, and larger memory capacity. Our study spans both sparse (mixture of experts) and dense transformer-based LLMs, revealing how system design choices affect Model FLOPS Utilization (MFU = Model flops per token $\times$ Observed tokens per sec/Peak flops of the hardware) and overall throughput. For the co-design study, we utilized an analytical performance modeling tool capable of predicting LLM runtime within 10
@@ 2.2.1. Model Components:
The MLP module includes three linear layers: FFup (feed-forward up projection), FFgate (feed-forward gating), and FFdown (feed-forward down projection). FFup and FFgate project the input from dimension $h$ to $4h$. FFgate’s output is activated non-linearly and element-wise multiplied with FFup’s output. FFdown then projects this result back to the original dimension. The MLP output is added to its input and normalized, serving as the input to the next decoder layer (Bambhaniya et al., 2024).
@@ 2.3. Systems and Their Parameters
Prior research (Yin et al., 2013) and recent advancements (Patel et al., 2025; Lightmatter, 2025; Systems, 2025) suggest the feasibility of a fully optical networks with co-packaged optics (CPO) for both HBD/LBD domains, offering high bandwidth, reduced power consumption, and cost-effectiveness. These networks can adopt 2D or 3D configurations, such as HyperX or PolarFly, allowing access to any GPU in two or three hops. A Full flat network use advanced CPO to equalize bandwidth between HBD and LBD, though communication outside HBD incurs higher latency due to multiple hops. Figure 4 shows a few potential Full flat topologies, with GPUs linked via single-hop connections within HBD, while those requiring multiple switches are part of LBD. FullFlat employs 2D/3D HyperX, Polar* or similar meshes, ensuring each node is $\leq$3 hops from any other at equal high bandwidth per hop. Adaptive routing mitigates “traffic storms” and reduces tail latency. FullFlat equalizes HBD and LBD (high and low bandwidth domains), allowing all traffic types on a single network without reconfiguration. This removes the scale-up/scale-out partitioning required in two-tier fabrics, simplifying scheduling and model mapping. Full flat refers to full bisection bandwidth over a folded high-radix, multi-stage optical mesh, not $N^{2}$ physical links. FullFlat offers adaptive routing, flexible workload placement, and unique parallelization opportunities. It is less sensitive to missing optimizations (compute-comm overlap, hardware collectives), with only a 5% performance gap across top 5000 configs (vs. 80% for two-tier), which should translate to less performance engineering effort (1). Additionally, it should also improve resiliency, robustness, and require less redundancy tax since any GPU can replace any faulty one anywhere in the network. While technical details of the Full flat network will likely be disclosed in a future paper, this study anticipates the development of such networks by 2026/2027.
@@
Table 3 outlines the system specifications used in our co-design analysis, informed by projections of future technological landscapes. The configuration FullFlat refers to the full flat system with equal scale-up and scale-out bandwidths. TwoTier-HBD64—128 represents upcoming two-tiered network with around $8\times$ difference between scale-up and scale-out bandwidths, experimented with HBD sizes of 64 and 128. TwoTier-HBD8 reflects current systems with an HBD size of 8. The GPUs used for FullFlat and TwoTier-HBD64—128 are of same computational power and storage, whereas GPUs used in TwoTier-HBD8 are weaker. Our co-design study examines LLM performance sensitivity to various system parameters in Table 3, providing insights into the system requirements for such models. We considered two-tiered networks with scale-up bandwidths of 450-3600 GB/s and scale-out bandwidths of 50-1600 GB/s, covering existing and potential future systems. The FLOPS FP8 and FP16 rows indicate floating-point operations per second for FP8 and FP16 precision, measured in petaflops. Tier 1 and Tier 2 Memory Bandwidth represent data transfer rates within the memory hierarchy, with Tier 1 denoting faster memory like HBM and Tier 2 being standard DDR/LPDDR memory. Memory Capacity reflects storage capacity in each tier. Scale-up and scale-out Bandwidth values describe intra-node/HBD and inter-node/LBD bandwidth capabilities. Latency indicates the minimum delay in data transfer for each network tier. Cluster Size denotes the number of nodes within each system. We anticipate systems of this scale will be available by 2026, and we analyze their performance on large-scale LLMs.
@@ 3. Co-design Study
HBDxx There are $x$ compute engines in the HBD domain
SUnn $n$ GB/s Bandwidth for the scale-up/HBD domain
SOmm $m$ GB/s Bandwidth for the scale-out/LBD domain
@@ 3.1. Strong Scaling with Cluster Size
By 2026, systems are anticipated to run the GPT-1.8T model at speeds 5050-$70\times$ faster (at 4K gpus, TwoTier-HBD8 vs others), assuming advanced networking infrastructures are available. However, increasing the number of GPUs may lead to performance degradation beyond a certain point due to overheads compared to the pure compute being done. This underscores the importance of strategically investing in advanced network architectures for large-scale deployments of LLMs.
@@ 3.2. Compute/Communication Overlap
High-performance interconnects like NVIDIA Mellanox InfiniBand and Intel Tofino switches (Corporation, 2020) support hardware-accelerated in-network collectives for scale-out domains, while NVSwitches offer accelerated collectives for scale-up domains. Nvidia’s SHARP FP8 technology allows CPUs and GPUs to offload collective tasks to the network chip (Moore, 2025; Graham et al., 2020) improving performance. The analytical model considers traffic volume, latency reduction, and GPU cycle savings (about 13%) due to the presence of a hardware-supported collective engine. Figure 5 (c) illustrates the performance slowdown from using software-based collectives, with $2\times$ traffic for all-reduce and $1.5\times$ for reduce-scatter and all-gather. The y-axis shows the percentage slowdown, with the number of GPUs varying from 64 (512 for GPT-29T) to 65,536 shown in the x-axis.
@@ 3.5. Sensitivity to Scale up (SU) Bandwidth
The most significant performance boost occurs when the SU bandwidth doubles from 450 GB/s to 900 GB/s. For GPT-1.8T with an HBD of 128, a $2\times$ increase in SU bandwidth results in a $1.6\times$ performance improvement. In contrast, with an HBD of 64, the improvement is $1.4\times$ because it relies on the slower SO network for some expert communications when using 8K GPUs. When the HBD is set to 128, experts can fully utilize the high-speed HBD/SU domain, leading to over a $2.62\times$ improvement from an $8\times$ increase in SU bandwidth. For GPT-29T, the performance is similar for both HBD64 and HBD128, as expert communication primarily depends on the SO domain in both scenarios. In this case, the $8\times$ increase in SU bandwidth provides a $1.9\times$ improvement.
@@ 3.6. Sensitivity to Scale-Out (SO) Bandwidth
Figure 5 (f) shows the impact of doubling scale-out (SO) bandwidth on GPT4 MOE models’ performance, with scale-up bandwidth fixed at 1,600 GB/s at 80% efficiency. We tested sensitivity for HBD sizes 64 and 128. For GPT-1.8T, in the HBD64 scenario, expert parallelization and communication occur through the scale-out domain, leading to consistent performance improvements with increased SO bandwidth ($1.36\times$ at 3.6TB/s). In the HBD128 case, since experts fit within HBD, improvements are limited to data and pipeline parallel communications with minimal impact (1%) on overall performance. For GPT-29T, in both HBD64 and HBD128 scenarios, most expert communications occur through the scale-out domain, resulting in steady performance gains with increased SO bandwidth ($1.35\times$ at 3.6TB/s). HBD128 shows slightly more improvements than HBD64 due to faster communication (i.e., all-to-all) within the HBD domain that the experts can leverage.
@@ 3.7. Sensitivity to Compute FLOPS
Figure 5 (g) illustrates the sensitivity of GPT4 MOE’s performance to varying FLOPS per GPU. The SU bandwidth is fixed at 1.6 TB/s, and the SO bandwidth is set at 200 GB/s for the TwoTier-HBD* configuration. The relative speedup compared to the baseline: TwoTier-HBD64 at 2.3 FP16 TFLOP/s per GPU. As FLOPS increase from 2.3 PF/s to 18.4 PF/s, throughput improves across all configurations: TwoTier-HBD64, TwoTier-HBD128, and FullFlat. For the GPT-1.8T model, TwoTier-HBD64 sees a $1.4\times$ boost, while TwoTier-HBD128 and FullFlat improve by $1.59\times$ and $1.66\times$, respectively. The performance gap between TwoTier-HBD64 and TwoTier-HBD128 ranges from 25% to 39%. For the GPT-29T model, TwoTier-HBD64 and TwoTier-HBD128 exhibit performance gains around $1.62\times$, whereas FullFlat achieves a $1.9\times$ gain with increased FLOPS.
@@ 3.8. Sensitivity to Memory Bandwidth
Figure 5 (h) shows the impact of varying HBM bandwidth from 3 TB/s to 48 TB/s across three configurations: TwoTier-HBD64, TwoTier-HBD128, and FullFlat, evaluated on 8,192 GPUs. For GPT-1.8T, increasing HBM bandwidth results in a $4.5\times$ performance improvement with a $16\times$ bandwidth increase. FullFlat consistently delivers the highest throughput, followed closely by TwoTier-HBD128, due to its capacity to accommodate most communications within the scale-up (HBD=128) domain for GPT-1.8T, resulting in less than a 1% performance gap between TwoTier-HBD128 and FullFlat.
For GPT-29T, increasing HBM bandwidth enhances performance across all configurations, achieving up to $3.2\times$ improvement. FullFlat consistently gives the best performance, while TwoTier-HBD64 and TwoTier-HBD128 exhibit similar performance because both HBD domains cannot accommodate all TP/ES/EP communications in the SU domain, necessitating substantial use of the slower scale-out domain.
@@ 3.9.1. Impact of Tier-1 Memory:
The benefit of co-design study is that it tells us exactly when to stop adding more resources in the system as it won’t bring provide any more ROI. This analysis indicates that having a 1.3TB capacity is a decent target for data centers for the next couple of years. Increasing HBM capacity boosts throughput significantly ($4.9\times$ for GPT-1.8T). Thus, optimizing HBM capacity alongside other system parameters is crucial for achieving optimal performance for LLMs.
@@ 3.10. Non MOE Dense LLM: GPT3-175B
Figure 7 shows the GPT-3 175B model’s performance sensitivity across different system setups. Throughput can be $10\times$ higher, with optimal scaling stabilizing at 32K GPUs. Performance is compute-bound until communication and other overheads become significant. Overlapping compute and communication is vital for smaller models, as neglecting this can cut performance by 43% at 16,384 nodes. Lack of hardware collectives may lead to a 29% drop. While HBD = 64 suffices for GPT-3, HBD = 128 is optimal for all models analyzed, given a scale-out bandwidth of 200 GB/s or more. This underscores the necessity for fast, efficient networks, hardware-accelerated collectives, and topology-aware communication libraries to effectively support LLM models in future data centers.
@@
Average: TwoTier-HBD8 19%, TwoTier-HBD64 6%, FullFlat 4%
@@
Table 6 presents the percentage of exposed communication and overhead times (recomputation, pipeline bubble, exposed offload times) for the three models discussed. MoE models exhibit significantly higher communication and overhead times compared to smaller dense models. The dense GPT3 model has higher arithmetic intensity and lower memory and network intensities compared to the MOE models. As a result, compared to the sparse MOE models, the dense model demonstrates a lower sensitivity to changes in scale-up and scale-out bandwidth. Sensitivity to HBM bandwidth increase is also lower (less than $3\times$ vs $4.5\times$ for the MOE model). However, flop sensitivity is higher than MOE counterparts ($3\times$ vs $2\times$).
@@
[…]
@@
[…]
@@
[…]